OpenStudy (brooke..help00):

@LOWKEY.S This is from a while back but I got a chance to redo it again. Please help. WILL MEDAL// FAN. Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (otherworldly):

They both have the same amount of solutions

OpenStudy (otherworldly):

isn't this 2.03

OpenStudy (brooke..help00):

How'd you know lol?

OpenStudy (lowkey.s):

This is known as an inequality. To solve the inequality you would bring the numbers out of the "ABSOLUTE VALUES". For each inequality you will have a negative and a positive solution. Ex: |5x+6|=41 .... that would result to being 5x+6=41 and 5x+6=-41. Then you will just solve for x on each side. You will get a solution set for both. Equation 1 and 2 have the same solution set. Each inequality gives you two solution sets .... a negative and a positive.

OpenStudy (lowkey.s):

your in FLVS right?

OpenStudy (lowkey.s):

cause i am too

OpenStudy (brooke..help00):

I am- lol. What grade?

OpenStudy (otherworldly):

me to

OpenStudy (brooke..help00):

I'm in 8th

OpenStudy (otherworldly):

im in 9th

OpenStudy (brooke..help00):

Oh- ok. So what is the correct answer?

OpenStudy (otherworldly):

^_^ i said it

OpenStudy (brooke..help00):

Nevermin- I got confused. Between how the second was worded lol

OpenStudy (otherworldly):

k